OpenStudy (airabell):

when graphed, which equation will yield the same maximum value as the graph of y=3x^2+6x+9 y=-3(x-1)^2-12 y=-3(x+1)^2-12 y=3(x-1)^2-12 y=3(x+1)^2+12

OpenStudy (irishboy123):

so this is an exercise in completing the square strikes me that the mental arithmetic will be simplified by noting that \(y=3x^2+6x+9 \implies \dfrac{y}{3} = x^2 + 2x + 3\) and working ot from there :-) much easier to square things etc BUT then putting the 3 back on the RHS so you can compare against your list and then you see the problem!! is there a typo in there??

OpenStudy (airabell):

thanks i got it from here. :)